# Insecurity Through Obscurity: Veiled Vulnerabilities in Closed-Source Contracts

### v1

@@ Appendix A Destroyer Inu Attack Analysis
Appendix B Historical Versions of MEV Bots
@@
Although smart contracts on Ethereum are generally immutable once deployed, their bytecode at a given address may change over time through the use of SELFDESTRUCT and CREATE2 opcodes (Wood et al., 2014). Specifically, SELFDESTRUCT removes the contract code from the state, and CREATE2 allows redeployment of a new contract at the same address using a fixed deployer address, salt, and modified initialization code. As a result, a smart contract may have different bytecode at different points in time (Ma et al., 2023).
@@
Figure 9. The distribution of the number of versions per MEV bot in our dataset.
[…]
@@
Loss estimation.   To approximate the maximum possible loss, we analyze the highest ERC-20 token balance held by each exploitable version of an MEV bot. Specifically, for each exploitable version, we identify the block range during which that version was active and observe the maximum ERC-20 token balance within that range. If the version is the latest one, we consider the range from its deployment up to block 22,000,000 (corresponding to March 8, 2025).
Consistent with the estimation for current loss, if a contract’s vulnerability is limited to a specific ERC-20 token, we consider only the potential loss associated with that token. If an MEV bot has multiple exploitable versions, we report the loss corresponding to the version with the highest observed ERC-20 balance. To simplify the estimation, we assume that all tokens were sold on March 8, 2025. This allows us to use the token prices on Binance (Binance, 2025) on that day to estimate the potential loss for seven major ERC-20 tokens: WETH, WBTC, USDC, USDT, DAI, UNI and LINK.
Appendix C Solidity Example of Checking Caller
@@ 18 POOL_INIT_CODE_HASH
23}
Figure 10. A Solidity implementation to check if the caller is a Uniswap V3 pool. It verifies whether a given pool address is valid by deriving its expected address using CREATE2.
[…]
A Solidity implementation to check if the caller is a Uniswap V3 pool. It verifies whether a given pool address is valid by deriving its expected address using CREATE2.

### v3

@@ Appendix B Case Study: the Destroyer Inu Attack
Appendix C Other Obfuscation Techniques
We identify four contract obfuscation techniques based on previous works [11, 81, 84, 82], summarized as:
  • Layout obfuscation typically aims to reduce information available to a human reader, e.g., by removing comments and renaming functions and variables [11, 84].
  • Data flow obfuscation obscures how data is processed and accessed within the smart contract [11, 84].
  • Control flow obfuscation alters the execution path of the smart contract to make its logical flow difficult to follow [11, 45], which is our focus.
  • Preventive transformations aim to hinder reverse engineering by exploiting weaknesses in current decompilers and deobfuscators [11]. E.g., contracts can self-destruct after execution to render their bytecode unavailable, preventing future inspection [82].
We empirically verified that the layout and data flow obfuscation do not interfere with existing analysis tools.
Two mature obfuscation tools are available for EVM smart contracts [82]: BOSC [81] and BiAn [84]. Since BOSC cannot guarantee that the obfuscated contracts remain deployable, it is not suitable for our setting. Following prior work [82], which also only used BiAn due to BOSC’s limitations, we adopt BiAn for our evaluation.
We collected vulnerable smart contracts from the Smart Contract Weakness Classification Registry [65] and applied BiAn to each contract to generate an obfuscated version. Specifically, we apply two types of obfuscation: layout obfuscation and data flow obfuscation. We did not apply control-flow obfuscation, since the current implementation in BiAn is non-functional [82, 76]. We then compile the original vulnerable contracts and the obfuscated ones, and apply Mythril to detect their vulnerabilities.
Interestingly, we found that for 65 of the vulnerable smart contracts, layout and data flow obfuscation do not prevent the vulnerabilities from being detected; instead, the obfuscation even introduces new “Assert Violation” issues [66]. For the remaining two contracts, we observed that Mythril could still detect issues in the obfuscated contracts at the source code level, but not at the bytecode level. We speculate this might be related to compiler optimizations. In summary, our findings indicate that existing analysis tools like Mythril can still work effectively on smart contracts with both layout and data flow obfuscation.
Appendix D Additional Details on Taint Tracking
At the bytecode level, Solidity’s complex data structures (e.g., mappings, arrays, structs) are compiled into memory offsets, calldata offsets, or storage slots. Accordingly, our taint tracking does not require field- or pointer-level semantics at the source level. Instead, we track taint at the granularity of offsets and slots: if a tainted value contributes to a memory offset or storage slot, the corresponding access is marked tainted. This offset- and slot-based tracking naturally captures fine-grained dependencies in high-level data structures.
Our taint analysis focuses on explicit data flows. We propagate taint only through data dependencies in the EVM (stack, memory, and storage) and do not model control dependencies (e.g., tainted branch conditions). This design choice ensures that our analysis captures whether adversarial calldata bytes directly influence the target address, function selector, or other arguments of a CALL instruction.
Appendix E Statistics on the Branch Table
We count the number of branches in each inserted branch table for the obfuscated contracts, and the results are shown in Fig. 9. It shows that 38% of the contracts have fewer than 100 branches, and about 70% have fewer than 400 branches, with a long tail extending to contracts containing over a thousand branches. Overall, the majority of contracts contain relatively few branches in their branch tables, and we find that a high number of branches does not necessarily imply greater complexity, since none of the inserted branch table caused a timeout during our evaluation.
Figure 9: Number of branches in the branch table of each deobfuscated smart contract.
Appendix F Evaluation of the Pruning Rule
To evaluate the pruning rule’s effectiveness, we conduct a comparative study by running skanf in full symbolic execution mode both without pruning and with a rule that limits the branch table to at most two accesses per execution. We focus on contracts in our datasets (see Sec. 4.1) that contain a branch table after deobfuscation.
Our evaluation shows that our pruning rule simplifies the state space in two cases, enabling skanf to detect vulnerabilities that would otherwise be missed. For the remaining contracts, the outcomes are identical with or without pruning. We attribute this to the large number of explored paths, where the accumulation of constraints during execution often makes the path unsatisfiable, leading the solver to discard it even without explicit pruning. Recall that we have exploited 162 obfuscated smart contracts (in Sec. 4.3), indicating that skanf is already effective in handling such cases, though further improvements in this direction remain an avenue for future work.
Appendix G Historical Versions of MEV Bots
@@
Although smart contracts on Ethereum are generally immutable once deployed, their bytecode at a given address may change over time through the use of SELFDESTRUCT and CREATE2 opcodes [75]. Specifically, SELFDESTRUCT removes the contract code from the state, and CREATE2 allows redeployment of a new contract at the same address using a fixed deployer address, salt, and modified initialization code. As a result, a smart contract may have different bytecode at different points in time [44].
@@
Figure 10: Distribution of the number of versions per MEV bot in our dataset.
@@
Loss estimation.  To approximate the maximum possible loss, we analyze the highest ERC-20 token balance held by each exploitable version of an MEV bot. Specifically, for each exploitable version, we identify the block range during which that version was active and observe the maximum ERC-20 token balance within that range. If the version is the latest one, we consider the range from its deployment up to block 22,635,000 (corresponding to June 5th, 2025).
Consistent with the estimation for current loss, if a contract’s vulnerability is limited to a specific ERC-20 token, we consider only the potential loss associated with that token. If an MEV bot has multiple exploitable versions, we report the loss corresponding to the version with the highest observed ERC-20 balance. To simplify the estimation, we assume that all tokens were sold on June 5th, 2025. This allows us to use the token prices on Binance [2] on that day to estimate the potential loss for seven major tokens: WETH, WBTC, USDC, USDT, DAI, UNI and LINK.
Appendix H Solidity Example of Checking Caller
@@ 16 POOL_INIT_CODE_HASH
20}
Figure 11: A Solidity implementation to check if the caller is a Uniswap V3 pool. It verifies whether a given pool address is valid by deriving its expected address using CREATE2.